OpenStudy (anonymous):

The solution set of the equation Cos(x-50)=-1/2; 0

OpenStudy (anonymous):

you can use the Trig Formula \[\cos(\alpha \pm \beta)=\cos \alpha \cos \beta \mp \sin \alpha \sin \beta\]